OpenStudy (anonymous):

in what base does 10=57 base 8 and 11=31 base 4? And why?

ganeshie8 (ganeshie8):

may i knw if they're two different questions or both are one q only

OpenStudy (anonymous):

They are actually two different questions.

Parth (parthkohli):

\[\rm 5\cdot 8 + 7 \cdot 1 = 47 \]

ganeshie8 (ganeshie8):

thank you :) Q1 : \((10)_x = (57)_8\) Q2 : \((11)_y = (31)_4\)

ganeshie8 (ganeshie8):

see parth's hint... you wil be able to figure out....

Parth (parthkohli):

\[\rm (10)_x = 57_8 \implies 1x^1 + 0x^0 = 57_8 \implies x = 57_8\]

Parth (parthkohli):

That's a ridiculous base. lol